A cube of glass of side 5cm and mass 300g has a cavity inside it. If the density of the glass is 2.5g/cm³, calculate the volume of cavity in the glass.
1
Expert's answer
2018-12-17T08:00:10-0500

The volume of emptiness is equal to the difference between the volume of the cube and the volume of glass:

V_cavity=V_cube-V_glass

Volume of cube:

V_cube=a^3

Volume of glass:

V_glass=m/ρ

V_cavity=a^3-m/ρ=125-300/2.5=5cm^3